OpenStudy (anonymous):

What is the solution to the system of equations? [-2x+2y-2=10_-3x+y-z=5_2x-3y+z=-1]

OpenStudy (mertsj):

Is there supposed to be a "z" in the first equation?

OpenStudy (anonymous):

Is That A Typo ?

OpenStudy (anonymous):

Oh...I see what he did...he used an underscore for a separator and had everything boxed inside brackets. Here's the untangled mess: -2x+2y-2=10 -3x+y-z=5 2x-3y+z=-1

OpenStudy (anonymous):

In That Case I Think This Website should help a bit http://www.algebra.com/algebra/homework/Human-and-algebraic-language/Human-and-algebraic-language.faq.question.266793.html (: